what is the volume of a sphere who's radius is
9 yards?

Answer by Theo(13342)

the formula for the volume of a sphere is:
V = 4/3*pi*r^3
if r = 9 yards, then the volume of the sphere is equal to:
V = 4/3 * pi * (9)^3 = 4/3 * pi * 729 = 4 * pi * 243 = 972 * pi.
pi is a constant that is equal to 3.141592654...
the ... indicates that this is a never ending decimal that has been rounded to 9 decimal places.